OpenStudy (anonymous):

Harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation. Part B: How many liters of apple juice are present in the jar? Show your work.

OpenStudy (anonymous):

@christiansousley

OpenStudy (anonymous):

You are given water = 1, so plugging in one for x, y=x+3x y=1+3 y=4

OpenStudy (anonymous):

75% represents unknown quantity of juice. 25% represents one liter of water. can you continue from here?

OpenStudy (anonymous):

Y=12?

OpenStudy (anonymous):

i think so?

OpenStudy (anonymous):

Haha okay